### 17 Jan 2024 v4

@@ Abstract
Tropical cyclones (TCs), driven by heat exchange between the air and sea, pose a substantial risk to many communities around the world. Accurate characterization of the subsurface ocean thermal response to TC passage is crucial for accurate TC intensity forecasts and for an understanding of the role that TCs play in the global climate system. However, that characterization is complicated by the high-noise ocean environment, correlations inherent in spatio-temporal data, relative scarcity of in situ observations, and the entanglement of the TC-induced signal with seasonal signals. We present a general methodological framework that addresses these difficulties, integrating existing techniques in seasonal mean field estimation, Gaussian process modeling, and nonparametric regression into an ANOVA decomposition model. Importantly, we improve upon past work by properly handling seasonality, providing rigorous uncertainty quantification, and treating time as a continuous variable, rather than producing estimates that are binned in time. This ANOVA model is estimated using in situ subsurface temperature profiles from the Argo fleet of autonomous floats through a multi-step procedure, which (1) characterizes the upper ocean seasonal shift during the TC season; (2) models the variability in the temperature observations; (3) fits a thin plate spline using the variability estimates to account for heteroskedasticity and correlation between the observations. This spline fit reveals the ocean thermal response to TC passage. Through this framework, we obtain new scientific insights into the interaction between TCs and the ocean on a global scale, including a three-dimensional characterization of the near-surface and subsurface cooling along the TC storm track and the mixing-induced subsurface warming on the track’s right side.
@@
Even after decades of research on TCs, a full understanding of the physics of the air–sea interface at high wind speeds is missing, which hinders our ability to simulate the structure and intensity of hurricane-strength storms (Emanuel, 2018). The principal aim of this paper is to provide statistical methodology to characterize temperature changes in the near-surface and subsurface ocean associated with the passage of a TC, given temperature observations from autonomous robotic devices called Argo floats (Jayne et al., 2017). In a naive sense, one may observe the TC-induced temperature change merely by comparing Argo temperature values, as a function of depth, recorded before and after the passage of a TC, as in Figure 1(b). However, the scarcity of such observations makes it nontrivial to obtain a more detailed characterization of the TC-induced signal. Here we develop methodology that enables such a characterization as a function of perpendicular distance from the TC track, time since TC passage, and subsurface depth. We call this the TC-induced signal (or “TC signal”) in a TC-centric coordinate system.
Challenges that remain in improving estimates of TC-induced ocean temperature changes include: (1) the high-noise environment of the ocean, coupled with the relative scarcity of in situ observations, which makes it difficult to observe statistically meaningful effects; (2) the need to disentangle the TC-induced signal from other, unrelated signals also present in the data, such as a seasonal warming effect that occurs in the ocean during the summer months; and (3) the highly correlated nature of the spatio-temporal Argo profiles, which necessitates careful estimation of the within- and between-pair covariances. In addressing each of these obstacles posed by the complex data model within which we work, we obtain an estimate of the TC-induced signal with a reasonable signal-to-noise ratio.
@@
For these seasonally adjusted temperature differences, we further detail a method for estimating their self-variances and cross-covariances, based upon the locally stationary Gaussian process model of Kuusela and Stein (2018). The reasons for fitting these covariances prior to producing the final fits are twofold: ocean dynamics conspire to produce additional heteroskedastic noise at large time lags; and pairs of profiles, especially those nearby in space and time, can be highly correlated. These covariance estimates are then used in a thin plate spline smoother to account for the effect of the correlated, heteroskedastic observations and to provide pointwise confidence intervals. Through these steps, we attain the chief scientific contributions of characterizing the ocean response to TC passage in the continuous time realm, rather than using a binned analysis as in Cheng et al. (2015), as well as properly accounting for the seasonal variation in ocean temperatures over the time scales of the paired profiles. Moreover, our estimation of within- and between-pair covariances, which were not accounted for in Cheng et al. (2015), admits more accurate fits and enables rigorous uncertainty quantification. These statistical procedures are unified in terms of an ANOVA decomposition, which is fit separately over a grid of twenty depth levels.
@@
Through this model, we more accurately characterize ocean cooling in the near-surface and subsurface waters, as well as the mixing induced warming in the subsurface on one side of the TC center. For a fixed depth, this characterization of the TC signal is performed by estimating a continuous function of cross-track distance and time since TC passage. We then repeat this estimation over a fine grid of depths from the ocean surface to 200 meters below the surface, using Argo profiles and TC track data from 2007 to 2018. In particular, the simultaneous estimation of the temperature change over these three axes (cross-track distance, time since TC passage, and depth) allows us to produce a three-dimensional characterization of the aforementioned scientific phenomena.
@@
The organization of the paper is as follows. Section 1 formally motivates, in oceanographic, meteorological, and climatological terms, the scientific problem of estimating the ocean thermal response to tropical cyclone passage. Section 2 introduces the tropical cyclone track data and the Argo profile data. Section 3 describes our methodology, where we begin by presenting the ANOVA-type data model and its implied data analysis pipeline. Data preprocessing is reviewed in Section 3.1, and the process for pairing profiles and projecting pairs onto TC tracks is detailed in Section 3.2. A local mean field model for capturing TC-unrelated seasonal effects is described in Section 3.3, followed by a Gaussian process model for estimating ocean variability in Section 3.4. In Section 3.5, we describe a version of the thin plate spline that allows for covariance reweighing. Section 4 describes the results of applying our methodology to the estimation of the global ocean thermal response to TCs. We conclude with a discussion of the results and future directions in Section 12.
@@
Appendix A provides a comprehensive review of the motivation and derivation for the thin plate spline variant used, and Appendix B details the leave-one-out cross-validation procedure used to select the regularization parameter. A full set of model fits is provided in the online supplement (Hu et al., 2020). In the spirit of reproducibility and to encourage the extension of these results, the code used to process the data and fit these models is freely available online at https://github.com/huisaddison/tc-ocean-methods.
@@ 1 Scientific context
As TCs intensify, their strengthening surface winds extract energy from the ocean surface and induce mixing of the upper ocean. This combination of processes decreases SST (creating a cold wake) and thus decreases the energy available for further TC intensification (Bender and Ginis, 2000; Shay, 2010). Also, the magnitude of mixing is proportional to the strength of the TC’s surface winds. Therefore, knowledge of subsurface ocean conditions can impact the accuracy of TC intensity forecasts (e.g., Mainelli et al., 2008). Such information, however, is difficult to obtain through remotely sensed, satellite-based observations, necessitating the introduction of in situ measurements, such as those provided by the Argo float program (Jayne et al., 2017).
@@
With more than two million profiles since the early 2000s, Argo floats provide unprecedented spatial and temporal coverage of the global ocean in the upper 2000 meters (Roemmich and Gilson, 2009; Riser et al., 2016; Jayne et al., 2017). Due to their autonomous nature, Argo floats are well-positioned to sample the ocean state in three dimensions before and after TC passage. Several previous studies have leveraged Argo profiles to describe TC-related changes in the upper ocean (Liu et al., 2007; Balaguru et al., 2015; Sun et al., 2012; Qu et al., 2014; Cheng et al., 2015; Lin et al., 2017; Steffen and Bourassa, 2018; Trenberth et al., 2018), focusing on changes in ocean temperature, salinity, or both, in order to analyze changes in ocean density. The closest work to ours is Cheng et al. (2015), who perform a binned analysis of temperature changes in time (compared to our continuous treatment of time) and do not account for a seasonal warming effect we discern in our analysis.
@@ 2.3 TC-centric coordinate system
Figure 2: Illustration of the coordinate systems with Hurricane Maria (2017). Figure 2(a) shows the path of the storm in red and the Argo profiles incidental to the storm in black. In Figure 2(b), we plot the paired profiles in the TC-centric coordinate system. Marginal histograms illustrate the density of the pairs along the time and cross-track axes.
@@ 3 Statistical methodology
The partitioning of the Argo profile database into TC and non-TC profile sets is described in Section 3.1.2. The pairing process and the construction of the observed temperature difference are detailed in Section 3.2. Section 3.3 explains the estimation of the seasonal effect based on spatio-temporal local regression modeling, and Section 3.4 describes a space-time Gaussian process model for the remaining ocean variability term. Finally, Section 3.5 leverages thin plate splines to produce a final fit of the TC signal $s(d,\tau)$. We depict the full framework as a data analysis pipeline in Figure 4.
@@ 3.2.1 Pairing process
We impose an additional technical restriction with respect to the lineage of a baseline profile, which is defined to be the set of all signal profiles associated with a distinct baseline profile and that baseline profile itself. We require that all profiles in the lineage of a baseline profile must be separated in time by at least 72 hours, sparsifying the lineage. To our knowledge, Cheng et al. (2015) imposes no such constraint. The choice of this separation is for a particular reason: we know that pairs of profiles close in time tend to be highly correlated, but our Gaussian process model (Kuusela and Stein, 2018) for estimating the between-pair correlations (a prerequisite for properly handling them in the final thin plate spline fit) may not provide reliable estimates in the rare cases where profiles occur more frequently than three days apart. Following a filter on minimum time separation within each lineage, we retain 16025 profile pairs across all five ocean regions: 2757 in the North Atlantic, 2897 in the East Pacific, 778 in the Indian Ocean, 2185 in the Southern Hemisphere, and 7409 in the West Pacific.
@@ 3.3 Seasonal mean field estimation
Using the non-TC profiles of Section 3.1.2, we estimate, separately for each pressure level $z=10,20,\dotsc,200$ and for the vertically averaged temperature values, a global mean field through a variant of the Roemmich–Gilson climatology. The Roemmich–Gilson climatology (Roemmich and Gilson, 2009) was among the first to take advantage of the spatial coverage of the Argo array to characterize the mean state and annual cycle of temperature in the global upper ocean. Specifically, they used a local linear regression model (Ridgway et al., 2002) for estimating the climatological mean temperature as a function of location $(x_{\mathrm{lon}},x_{\mathrm{lat}})$, pressure level $z$, and day of year $t$. This model is written down explicitly in Kuusela and Stein (2018), and our model, which differs marginally from that of Roemmich and Gilson (2009) in that we do not weigh observations based on distance from the center point $(x_{\mathrm{lon}}^{*},x_{\mathrm{lat}}^{*})$, nor do we use data from pressure levels $z^{\prime}\neq z$ to form estimates at pressure level $z$, is elucidated in Equation (6).
@@ 3.4.1 Computational remarks
Denote by $\boldsymbol{y}$ the full set of mean-adjusted temperature differences, ordered such that profiles in the same lineage are contiguous. Its covariance, $\operatorname{\mathrm{cov}}(\boldsymbol{y})$, is a block diagonal matrix, with dense blocks corresponding to the lineages of distinct baseline profiles. This is important for two reasons. First, a block diagonal matrix admits an efficient inverse; namely, the inverse of a block diagonal matrix may be obtained by inverting the blocks individually. Each of these blocks is no more than $6$ entries wide, due to the separation constraint placed on profiles in Section 3.2. Second, the sparse structure of the block diagonal matrix may be exploited for efficient linear algebra operations (Virtanen et al., 2020) necessary to construct the thin plate splines described in the sequel.
3.5 Thin plate splines
Having learned a model for the $(a(\boldsymbol{x},t_{\text{signal{}}})-a(\boldsymbol{x},t_{\text{baseline{}}}))$ term of (11) through the Gaussian process model of Section 3.4, we find ourselves properly positioned to estimate the TC-induced signal $s(d,\tau)$. This we do with a thin plate spline, a flexible nonparametric technique for fitting a surface in two dimensions, subject to a curvature penalty and observation weights. Specifically, we smooth the seasonally adjusted temperature differences using a fixed-knot thin plate spline smoother (Duchon, 1977; Wahba, 1980, 1990; Green and Silverman, 1994; Nychka, 2000; Wood, 2017), in which we use the covariances estimated in Section 3.4 to account for heteroskedasticity and cross-observation dependence (Rao, 1973; Draper and Smith, 1998; Nychka, 2000; Ruppert et al., 2003). As with the local mean field model of Section 3.3 and the Gaussian process model of Section 3.4, this procedure is conducted separately at each pressure level $z\in\{10,20,\dotsc,200\}$ and for the vertically averaged profiles.
@@ 3.5 Thin plate splines
The weighing of observations by the covariance between temperature differences is also noteworthy. From a data model standpoint, it is known that we are using observations with drastically different noise levels $a(x_{\mathrm{lon}},x_{\mathrm{lat}},t_{\text{baseline{}}})$, $a(x_{\mathrm{lon}},x_{\mathrm{lat}},t_{\text{signal{}}})$, due to the effect of time and location. Since we have estimated these noise levels in Section 3.4, the covariance-reweighing allows us to directly incorporate this information when estimating $s(d,\tau)$ in the reduced model (11). From a practical standpoint, the covariance-reweighing has a real effect in improving the signal-to-noise ratio, as shown in Figure 7 of Section 4.2. The full details of the thin plate spline smoother, including its motivation and derivation, are given in Appendix A.
@@
The fixed-knot covariance-reweighed thin plate spline problem is posed as follows. We seek coefficients $\boldsymbol{\delta},\boldsymbol{\beta}$ such that the following objective is minimized:

### 14 Mar 2024 v5

@@ Abstract
Tropical cyclones (TCs), driven by heat exchange between the air and sea, pose a substantial risk to many communities around the world. Accurate characterization of the subsurface ocean thermal response to TC passage is crucial for accurate TC intensity forecasts and for an understanding of the role that TCs play in the global climate system. However, that characterization is complicated by the high-noise ocean environment, correlations inherent in spatio-temporal data, relative scarcity of in situ observations, and the entanglement of the TC-induced signal with seasonal signals. We present a general methodological framework that addresses these difficulties, integrating existing techniques in seasonal mean field estimation, Gaussian process modeling, and nonparametric regression into an ANOVA decomposition model. Importantly, we improve upon past work by properly handling seasonality, providing rigorous uncertainty quantification, and treating time as a continuous variable, rather than producing estimates that are binned in time. This ANOVA model is estimated using in situ subsurface temperature profiles from the Argo fleet of autonomous floats through a multi-step procedure, which (1) characterizes the upper ocean seasonal shift during the TC season; (2) models the variability in the temperature observations; (3) fits a thin-plate spline using the variability estimates to account for heteroskedasticity and correlation between the observations. This spline fit reveals the ocean thermal response to TC passage. Through this framework, we obtain new scientific insights into the interaction between TCs and the ocean on a global scale, including a three-dimensional characterization of the near-surface and subsurface cooling along the TC storm track and the mixing-induced subsurface warming on the track’s right side.
@@
Even after decades of research on TCs, a full understanding of the physics of the air–sea interface at high wind speeds is missing, which hinders our ability to simulate the structure and intensity of hurricane-strength storms (Emanuel, 2018). The principal aim of this paper is to provide statistical methodology to characterize temperature changes in the near-surface and subsurface ocean associated with the passage of a TC, given temperature observations from autonomous robotic devices called Argo floats (Jayne et al., 2017). In a naive sense, one may observe the TC-induced temperature change merely by comparing Argo temperature values, as a function of depth, recorded before and after the passage of a TC, as in Figure 1(b). However, the scarcity of such observations makes it nontrivial to obtain a more detailed characterization of the TC-induced signal. Here we develop methodology that enables such a characterization as a function of perpendicular distance [^1] Measured in terms of angular distance to the nearest point on the TC track; alternately referred to as “cross-track distance” and “cross-track angle”. from the TC track, time since TC passage, and subsurface depth. We call this the TC-induced signal (or “TC signal”) in a TC-centric coordinate system.
Challenges that remain in improving estimates of TC-induced ocean temperature changes include: (1) the high-noise environment of the ocean, coupled with the relative scarcity of in situ observations, which makes it difficult to observe statistically meaningful effects; (2) the need to disentangle the TC-induced signal from other, unrelated signals also present in the data, such as a seasonal warming effect that occurs in the ocean during the summer months; and (3) the highly correlated nature of the spatio-temporal Argo profiles, which necessitates careful estimation of the covariance between profiles located close to each other in space and time. In addressing each of these obstacles posed by the complex data model within which we work, we obtain an estimate of the TC-induced signal with a reasonable signal-to-noise ratio.
@@
For these seasonally adjusted temperature differences, we further detail a method for estimating their self-variances and cross-covariances, based upon the locally stationary Gaussian process model of Kuusela and Stein (2018). The reasons for fitting these covariances prior to producing the final fits are twofold: ocean dynamics conspire to produce additional heteroskedastic noise at large time lags; and pairs of profiles, especially those nearby in space and time, can be highly correlated. These covariance estimates are then used in a thin-plate spline smoother to account for the effect of the correlated, heteroskedastic observations and to provide pointwise confidence intervals. Through these steps, we attain the chief scientific contributions of characterizing the ocean response to TC passage in the continuous time realm, rather than using a binned analysis as in Cheng et al. (2015), as well as properly accounting for the seasonal variation in ocean temperatures over the time scales of the paired profiles. Moreover, our estimation of covariances between profiles co-located in space and time, which were not accounted for in Cheng et al. (2015), admits more accurate fits and enables rigorous uncertainty quantification. These statistical procedures are unified in terms of an ANOVA decomposition, which is fit separately over a grid of twenty depth levels.
@@
Through this model, we more accurately characterize ocean cooling in the near-surface and subsurface waters, as well as the mixing induced warming in the subsurface on one side of the TC center. For a fixed depth, this characterization of the TC signal is performed by estimating a continuous function of cross-track distance and time since TC passage. We then repeat this estimation over a fine grid of depths from the ocean surface to 200 meters below the surface, using Argo profiles and TC track data from 2007 to 2018. In particular, since our estimates of the temperature change vary continuously over these three axes (cross-track distance, time since TC passage, and depth), we are able to produce a three-dimensional characterization of the aforementioned scientific phenomena.
@@
The organization of the paper is as follows. Section 1 formally motivates, in oceanographic, meteorological, and climatological terms, the scientific problem of estimating the ocean thermal response to tropical cyclone passage. Section 2 introduces the tropical cyclone track data and the Argo profile data. Section 3 describes our methodology, where we begin by presenting the ANOVA-type data model and its implied data analysis pipeline. Data preprocessing is reviewed in Section 3.1, and the process for pairing profiles and projecting pairs onto TC tracks is detailed in Section 3.2. A local mean field model for capturing TC-unrelated seasonal effects is described in Section 3.3, followed by a Gaussian process model for estimating ocean variability in Section 3.4. In Section 3.5, we describe a version of the thin-plate spline that allows for covariance reweighting. Section 4 describes the results of applying our methodology to the estimation of the global ocean thermal response to TCs. We conclude with a discussion of the results and future directions in Section 12.
@@
Appendix A provides a comprehensive review of the motivation and derivation for the thin-plate spline variant used, and Appendix B details the leave-one-out cross-validation procedure used to select the regularization parameter. A full set of model fits is provided in the online supplement (Hu et al., 2020). In the spirit of reproducibility and to encourage the extension of these results, the code used to process the data and fit these models is freely available online at https://github.com/huisaddison/tc-ocean-methods.
@@ 1 Scientific context
As TCs intensify, their strengthening surface winds extract energy from the ocean surface and induce mixing of the upper ocean. This combination of processes decreases SST (creating a cold wake) and thus decreases the energy available for further TC intensification (Bender and Ginis, 2000; Shay, 2010). The magnitude of mixing depends on the strength of the TC’s surface winds, its size, and translation speed, as well as background ocean conditions. Therefore, knowledge of subsurface ocean conditions can impact the accuracy of TC intensity forecasts (e.g., Mainelli et al., 2008). Such information, however, is difficult to obtain through remotely sensed, satellite-based observations, necessitating the introduction of in situ measurements, such as those provided by the Argo float program (Jayne et al., 2017).
@@
With more than two million profiles since the early 2000s, Argo floats provide unprecedented spatial and temporal coverage of the global ocean in the upper 2000 meters (Roemmich and Gilson, 2009; Riser et al., 2016; Jayne et al., 2017). Due to their autonomous nature, Argo floats are well-positioned to sample the ocean state in three dimensions before and after TC passage. Several previous studies have leveraged Argo profiles to describe TC-related changes in the upper ocean (Liu et al., 2007; Balaguru et al., 2015; Sun et al., 2012; Qu et al., 2014; Cheng et al., 2015; Lin et al., 2017; Steffen and Bourassa, 2018; Trenberth et al., 2018), focusing on changes in ocean temperature, salinity, or both, in order to analyze changes in ocean density. The closest work to ours is Cheng et al. (2015), who perform a binned analysis of temperature changes in time (compared to our continuous treatment of time) and do not account for a seasonal warming effect we discern in our analysis. An earlier paired analysis in the north Pacific was performed by Park et al. (2011).
@@ 2.3 TC-centric coordinate system
Figure 2: Illustration of the coordinate systems with Hurricane Maria (2017). Figure 2(a), using the latitude-longitude coordinate system, shows the path of the storm in red and the location of the Argo profile pairs straddling the storm passage in black. In Figure 2(b), we plot those same pairs of profiles in the TC-centric coordinate system. It is possible for one black point in the latitude-longitude coordinate system to correspond to multiple observation in the TC-centric coordinate system, if a single profile occurring before TC passage is paired with multiple profiles occurring after TC passage.
@@ 3 Statistical methodology
The partitioning of the Argo profile database into TC and non-TC profile sets is described in Section 3.1.2. The pairing process and the construction of the observed temperature difference are detailed in Section 3.2. Section 3.3 explains the estimation of the seasonal effect based on spatio-temporal local regression modeling, and Section 3.4 describes a space-time Gaussian process model for the remaining ocean variability term. Finally, Section 3.5 leverages thin-plate splines to produce a final fit of the TC signal $s(d,\tau)$. We depict the full framework as a data analysis pipeline in Figure 4.
@@ 3.2.1 Pairing process
We impose an additional technical restriction with respect to the lineage of a baseline profile, which is defined to be the set of all signal profiles associated with a distinct baseline profile and that baseline profile itself. We require that all profiles in the lineage of a baseline profile must be separated in time by at least 72 hours, sparsifying the lineage. To our knowledge, Cheng et al. (2015) imposes no such constraint. The choice of this separation is for a particular reason: we know that pairs of profiles close in time tend to be highly correlated, but our Gaussian process model (Kuusela and Stein, 2018) for estimating the correlations between pairs of profiles that are close in space and time (a prerequisite for properly handling them in the final thin-plate spline fit) may not provide reliable estimates in the rare cases where profiles occur more frequently than three days apart. Following a filter on minimum time separation within each lineage, we retain 16025 profile pairs across all five ocean regions: 2757 in the North Atlantic, 2897 in the East Pacific, 778 in the Indian Ocean, 2185 in the Southern Hemisphere, and 7409 in the West Pacific.
@@ 3.3 Seasonal mean field estimation
Using the non-TC profiles of Section 3.1.2, we estimate, separately for each pressure level $z=10,20,\dotsc,200$ and for the vertically averaged temperature values, a global mean field through a variant of the Roemmich–Gilson climatology. The Roemmich–Gilson climatology (Roemmich and Gilson, 2009) was among the first to take advantage of the spatial coverage of the Argo array to characterize the mean state and annual cycle of temperature in the global upper ocean. Specifically, they used a local linear regression model (Ridgway et al., 2002) for estimating the climatological mean temperature as a function of location $(x_{\mathrm{lon}},x_{\mathrm{lat}})$, pressure level $z$, and day of year $t$. This model is written down explicitly in Kuusela and Stein (2018), and our model, which differs marginally from that of Roemmich and Gilson (2009) in that we do not weight observations based on distance from the center point $(x_{\mathrm{lon}}^{*},x_{\mathrm{lat}}^{*})$, nor do we use data from pressure levels $z^{\prime}\neq z$ to form estimates at pressure level $z$, is elucidated in Equation (6).
@@ 3.4.1 Computational remarks
Denote by $\boldsymbol{y}$ the full set of mean-adjusted temperature differences, ordered such that profiles in the same lineage are contiguous. Its covariance, $\operatorname{\mathrm{cov}}(\boldsymbol{y})$, is a block diagonal matrix, with dense blocks corresponding to the lineages of distinct baseline profiles. This is important for two reasons. First, a block diagonal matrix admits an efficient inverse; namely, the inverse of a block diagonal matrix may be obtained by inverting the blocks individually. Each of these blocks is no more than $6$ entries wide, due to the separation constraint placed on profiles in Section 3.2. Second, the sparse structure of the block diagonal matrix may be exploited for efficient linear algebra operations (Virtanen et al., 2020) necessary to construct the thin-plate splines described in the sequel.
3.5 Thin-plate splines
Having learned a model for the $(a(\boldsymbol{x},t_{\text{signal{}}})-a(\boldsymbol{x},t_{\text{baseline{}}}))$ term of (11) through the Gaussian process model of Section 3.4, we find ourselves properly positioned to estimate the TC-induced signal $s(d,\tau)$. This we do with a thin-plate spline, a flexible nonparametric technique for fitting a surface in two dimensions, subject to a curvature penalty and observation weights. Specifically, we smooth the seasonally adjusted temperature differences using a fixed-knot thin-plate spline smoother (Duchon, 1977; Wahba, 1980, 1990; Green and Silverman, 1994; Nychka, 2000; Wood, 2017), in which we use the covariances estimated in Section 3.4 to account for heteroskedasticity and cross-observation dependence (Rao, 1973; Draper and Smith, 1998; Nychka, 2000; Ruppert et al., 2003). As with the local mean field model of Section 3.3 and the Gaussian process model of Section 3.4, this procedure is conducted separately at each pressure level $z\in\{10,20,\dotsc,200\}$ and for the vertically averaged profiles.
@@ 3.5 Thin-plate splines
The weighting of observations by the covariance between temperature differences is also noteworthy. From a data model standpoint, it is known that we are using observations with drastically different noise levels $a(x_{\mathrm{lon}},x_{\mathrm{lat}},t_{\text{baseline{}}})$, $a(x_{\mathrm{lon}},x_{\mathrm{lat}},t_{\text{signal{}}})$, due to the effect of time and location. Since we have estimated these noise levels in Section 3.4, the covariance-reweighting allows us to directly incorporate this information when estimating $s(d,\tau)$ in the reduced model (11). From a practical standpoint, the covariance-reweighting has a real effect in improving the signal-to-noise ratio, as shown in Figure 7 of Section 4.2. The full details of the thin-plate spline smoother, including its motivation and derivation, are given in Appendix A.
@@
The fixed-knot covariance-reweighted thin-plate spline problem is posed as follows. We seek coefficients $\boldsymbol{\delta},\boldsymbol{\beta}$ such that the following objective is minimized: